predict the concentration of water if you want to synthesize the polyester with 200 degree of polymerization. Keq = 2.0, Co = 1.0
Answer #1

According to Carrother's equation:

The number-average degree of polymerization (Xn) = 1/(1-p)

Where 'p' is the degree of polymerization and Xn = 200

Now, 200 = 1/(1-p)

i.e. 1-p = 1/200 = 0.005

Therefore, p = 1-0.005 = 0.995

Now, C = C0 * (1-p)

i.e. C = 1 * 0.005

Therefore, the concentration of water (C) = 0.005 M